# APOLLO Blender: A Robotics Library for Visualization and Animation in Blender

### v1

@@ I Introduction
In this work, we present APOLLO Blender, a lightweight software library that lowers this barrier by providing robotics-focused scripting tools within Blender. Our goal is not to replicate the full functionality of physics simulators or game engines, but rather to make it straightforward for researchers to create static images, animations, and schematic illustrations that are both visually appealing and robotics-aware. The library centers on three key features: importing robots and environments from standardized descriptions such as URDF, keyframing robot configurations and appearance attributes, and generating common 3D primitives such as lines and cubes to support schematics. Together, these capabilities streamline the visualization workflow, reducing the time and expertise required to produce publication-ready results.
@@ III-B Robot Importing
Robots and environments can be imported directly into Blender using standardized URDF descriptions. This process automatically loads both the geometric meshes that define the robot’s appearance and the kinematic structure that governs its configuration. By relying on URDF, researchers can reuse existing specifications without modification, ensuring compatibility with a wide range of commonly used platforms.
A typical script to spawn a robot requires only a few lines of code. The example below demonstrates how a URDF can be loaded and visualized:

### v2

@@ I Introduction
In this work, we present APOLLO Blender, a lightweight software library that lowers this barrier by providing robotics-focused scripting tools within Blender. Our goal is not to replicate the full functionality of physics simulators or game engines, but rather to make it straightforward for researchers to create static images, animations, and schematic illustrations that are both visually appealing and robotics-aware. The library centers on three key features: importing robots and environments from standardized descriptions, keyframing robot configurations and appearance attributes, and generating common 3D primitives such as lines and cubes to support schematics. Together, these capabilities streamline the visualization workflow, reducing the time and expertise required to produce publication-ready results.
@@ III-B Robot Importing
Robots and environments can be imported directly into Blender using a Universal Robot Description Directory (URDD) [klein2025urdd]. A URDD can be constructed automatically from a standard URDF. This process automatically loads both the geometric meshes that define the robot’s appearance and the kinematic structure that governs its configuration. By relying on URDD, researchers can reuse existing specifications without modification, ensuring compatibility with a wide range of commonly used platforms.
A typical script to spawn a robot requires only a few lines of code. The example below demonstrates how a URDD can be loaded and visualized: